# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 1.1. Hilbert bimodules
First, note that $\overline{\psi(B)C}=\psi(B)C$ by Cohen’s factorisation (see [37, Proposition 2.33]). Let $T:{}_{\phi}B\boxtimes{}_{\psi}C\to{}_{\psi\circ\phi}\psi(B)C$ be the continuous linear map given by $T(b\boxtimes c)=\psi(b)c$ for any $b\in B$ and $c\in C$. We claim that $T$ is a bimodule isomorphism. Using the definition of the inner product, a standard check shows that $T$ is a well-defined bimodule map. Taking $(\eta_{\lambda})_{\lambda\in\Lambda}$ to be an approximate unit for $B$ let $S:{}_{\psi\circ\phi}\psi(B)C\to{}_{\phi}B\boxtimes{}_{\psi}C$ be given by $S(c)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes c$ for any $c\in C$. The map $S$ is well-defined precisely because $\psi$ is extendible. Indeed, for $\lambda,\mu\in\Lambda$
@@
Therefore, $S$ is well defined if and only if $\psi(\eta_{\lambda}-\eta_{\mu})c$ converges to 00 for any $c\in C$, which is equivalent to saying that $\psi(\eta_{\lambda})$ converges strictly in $\mathcal{M}(C)$. Hence $S$ is well-defined. Moreover, it can be seen that $S$ is the adjoint of $T$ and hence continuous. Now for any $b\in B$ and $c\in C$,
@@
Let $A$ be a $*{}^{*}$-algebra. A Hilbert $A$-$B$-bimodule $X$ is called non-degenerate if $XB$ is dense in $X$ and $AX$ is dense in $X$.
@@
Note that if $X$ is a right-Hilbert $B$-module then $XB$ is always dense in $X$ by the argument in [25, Proposition 2.6]. Therefore, a Hilbert $A$-$B$-bimodule $E$ may only fail to be non-degenerate if $AX$ is not dense in $X$.
We end this subsection by showing that if $A$ is a $\mathrm{C}^{*}$-algebra and $E$ is a non-degenerate Hilbert $A$-bimodule, then we can obtain a well-defined action of $\mathcal{M}(A)$ on $E$.
Let $(\eta_{\lambda})_{\lambda\in\Lambda}$ be an approximate unit of $A$ and let $L_{E}:E\to A\boxtimes E$ and $R_{E}:E\to E\boxtimes A$ be the $A$-bimodule maps given by $L_{E}(x)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes x$ and $R_{E}(x)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}$ for all $x\in E$. Note that, since $E$ is non-degenerate, $L_{E}$ and $R_{E}$ are unitary bimodule isomorphisms. This follows similarly to the proof of Proposition 1.7. The inverses of $L_{E}$ and $R_{E}$ are given by their adjoints $L_{E}^{-1}(a\boxtimes x)=a\rhd x$ and $R_{E}^{-1}(x\boxtimes a)=x\lhd a$, for all $a\in A$ and $x\in E$ respectively.
@@
Let $A$ be a $\mathrm{C}^{*}$-algebra and $E$ be a non-degenerate Hilbert $A$-bimodule. Then one may extend the left and right actions of $A$ on $E$ to left and right actions of $\mathcal{M}(A)$ on $E$. These extended actions equip $E$ with the structure of a Hilbert $\mathcal{M}(A)$-bimodule.
@@
(1.1) $v\rhd x=L_{E}^{-1}(v\rhd L_{E}(x))\quad\mathrm{and}\quad x\lhd v=R_{E}^{-1}(R_{E}(x)\lhd v).$
We claim that these formulae define left and right actions of $\mathcal{M}(A)$ on $E$. Moreover, it is clear that (1.1) restricted to $A$ coincides with the $A$-bimodule structure of $E$.
First, note that the left action of $\mathcal{M}(A)$ on $A\boxtimes E$ is given by left multiplication on $A$. Similarly, the right action of $\mathcal{M}(A)$ on $E\boxtimes A$ is given by right multiplication on $A$. Since $L_{E},R_{E}$, and their inverses are bimodule maps, it is straightforward to see that the formulae in (1.1) define left and right actions of $\mathcal{M}(A)$ on $E$. To see that $E$ with its right $A$-valued inner product is a Hilbert $\mathcal{M}(A)$-bimodule it suffices to check that the right $\mathcal{M}(A)$-action commutes with the inner product and the left $\mathcal{M}(A)$-action consists of adjointable operators. First, for $x,y\in E$ and $v\in\mathcal{M}(A)$ we have that
@@
so the right $\mathcal{M}(A)$-action commutes with the right inner product. Moreover, the operator of left multiplication by $v\in\mathcal{M}(A)$ has as an adjoint the operator of left multiplication by $v^{*}$. Indeed left multiplication by $v^{*}$ is the adjoint of the operator of left multiplication by $v$ on the Hilbert $A$-module $A$ so for $x,y\in E$
@@ 1.2. Correspondences
A reformulation of the theory of Hilbert bimodules is the language of $*{}^{*}$-correspondences.
@@
Let $A,B$ be $\mathrm{C}^{*}$-algebras. An $A$-$B$ correspondence is a $*{}^{*}$-homomorphism $\phi:A\to\mathcal{L}(X_{B})$ into the adjointable operators of a Hilbert $B$-module $X_{B}$. We denote the collection of $A$-$B$ correspondences by $\mathrm{Corr}(A,B)$.
@@
Note that any $A$-$B$ correspondence induces a right-Hilbert $A$-$B$-bimodule and vice versa. If $\phi:A\to\mathcal{L}(X_{B})$ is an $A$-$B$ correspondence, then $X_{B}$ becomes a right-Hilbert $A$-$B$-bimodule, with the left action given by $\phi$. We will often denote this bimodule by ${}_{\phi}X$, forgetting the right $B$-action. Conversely, given a right-Hilbert $A$-$B$-bimodule $X$, the left action by $A$ induces an $A$-$B$ correspondence. Therefore, we will freely flip between the two pictures.
One may compose correspondences through the tensor product of bimodules. Precisely, let $X$ be a Hilbert $A$-$B$ bimodule inducing $\phi\in\mathrm{Corr}(A,B)$, and let $E$ be a Hilbert $B$-$C$ bimodule inducing $\psi\in\mathrm{Corr}(B,C)$. Then $X\boxtimes E$ gives a Hilbert $A$-$C$ bimodule which induces an element in $\mathrm{Corr}(A,C)$ denoted by $\psi\circ\phi$. Although the theories of bimodules and correspondences are equivalent, we sometimes choose to work with correspondences as the composition resembles composition of $*{}^{*}$-homomorphisms between $\mathrm{C}^{*}$-algebras in a covariant manner.
@@ 1.3. $\mathrm{C}^{*}$-tensor categories
Throughout this section we will assume that the reader is familiar with the standard language of category theory. For a category $\mathcal{C}$ we will use capital letters e.g. $X,Y,$ and $Z$ to denote objects of the category. The space of morphisms between two objects $X,Y\in\mathcal{C}$ will be denoted by $\mathrm{Hom}(X,Y)$. All categories in this section will be $\mathbb{C}$-linear, that is and additive category such that its space of morphisms $\mathrm{Hom}(X,Y)$ between any two objects $X,Y$ is a $\mathbb{C}$-vector space and the composition of morphisms yields a bilinear map $\mathrm{Hom}(X,Y)\times\mathrm{Hom}(Y,Z)\rightarrow\mathrm{Hom}(X,Z)$ for any $X,Y,Z\in\mathcal{C}$. We start by defining $\mathrm{C}^{*}$-categories.
@@
(see [19]) A $\mathbb{C}$-linear category $\mathcal{C}$ is said to be a $\rm{C}^{*}$-category if it is equipped with conjugate linear mappings :*${}^{*}\!:\mathrm{Hom}(X,Y)\rightarrow\mathrm{Hom}(Y,X)$ for every $X,Y\in\mathcal{C}$ such that
  1. (i)
    $\phi^{**}=\phi$ for all $\phi\in\mathrm{Hom}(X,Y);$
  2. (ii)
    $(\phi\circ\psi)^{*}=\psi^{*}\circ\phi^{*}$ for all $\psi\in\mathrm{Hom}(X,Y),\phi\in\mathrm{Hom}(Y,Z)$;
  3. (iii)
    The function $\|\cdot\|:\mathrm{Hom}(X,Y)\rightarrow[0,\infty]$ given by
    $\|\phi\|^{2}=\sup\{\lambda>0:\phi^{*}\circ\phi-\lambda\mathrm{id}_{X}\mbox{ isnot invertible}\}$
    is a complete norm on $\mathrm{Hom}(X,Y)$;
  4. (iv)
    $\|\phi\circ\psi\|\leq\|\phi\|\|\psi\|$ for all $\psi\in\mathrm{Hom}(X,Y),\phi\in\mathrm{Hom}(Y,Z)$;
  5. (v)
    $\|\phi^{*}\circ\phi\|=\|\phi\|^{2}$ for all $\phi\in\mathrm{Hom}(X,Y)$;
  6. (vi)
    For all $\phi\in\mathrm{Hom}(X,Y)$, $\phi^{*}\circ\phi$ is a positive element of the $\rm{C}^{*}$-algebra $\mathrm{Hom}(X,X)$.
Let $\mathcal{C}$ and $\mathcal{D}$ be $\mathrm{C}^{*}$-categories. A functor $F:\mathcal{C}\rightarrow D$ is called a $\mathrm{C}^{*}$-functor if the induced mappings $\mathrm{Hom}(X,Y)\mapsto\mathrm{Hom}(F(X),F(Y))$ are $\mathbb{C}$-linear and $*{}^{*}$-preserving. A natural transformation $\nu:F\rightarrow G$ between $\mathrm{C}^{*}$-functors is called an isometry if $\nu_{X}^{*}\nu_{X}=\mathrm{id}_{F(X)}$ for all $X\in\mathcal{C}$. Moreover, $\nu$ is called a unitary if it is a surjective isometry in which case $\nu_{X}\nu_{X}^{*}=\mathrm{id}_{G(x)}$ for all $X\in\mathcal{C}$. We are interested in $*{}^{*}$-categories that admit tensor product structures.
Definition 1.15.
(see for example [23]) A $\rm{C}^{*}$-tensor category is a $\rm{C}^{*}$-category $\mathcal{C}$ together with a $\mathbb{C}$-linear bifunctor $-\otimes-:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$, a distinguished object $1_{\mathcal{C}}\in\mathcal{C}$ and unitary natural isomorphisms
@@
[…]
and $\mathrm{Hom}(1_{\mathcal{C}},1_{\mathcal{C}})\cong\mathbb{C}$. Moreover, $\mathcal{C}$ is said to be semisimple if there exists a collection of objects $X_{i}\in\mathcal{C}$ with $\mathrm{Hom}(X_{i},X_{j})\cong\delta_{ij}\mathbb{C}$ and any object $X\in\mathcal{C}$ can be decomposed uniquely as a finite direct sum $X\cong\bigoplus_{i}X_{i}$.
We call the structure morphism $\alpha$ as in Definition 1.15 associated to a $*{}^{*}$-tensor category its associator and the maps $\lambda$ and $\rho$ the unitors. We call an object $X\in\mathcal{C}$ such that $\mathrm{Hom}(X,X)\cong\mathbb{C}$ irreducible or simple. For a $*{}^{*}$-tensor category $\mathcal{C}$, we will denote by $\mathrm{Irr}(\mathcal{C})$ a collection of isomorphism class representatives for simple objects in $\mathcal{C}$.
@@
Let $\Gamma$ be a countable, discrete group (or more generally a countable, discrete monoid). We denote by $\mathrm{Hilb}(\Gamma)$ the semisimple $*{}^{*}$-tensor category whose objects are finite-dimensional $\Gamma$-graded Hilbert spaces, i.e. finite-dimensional Hilbert spaces $\mathcal{H}$ endowed with a decomposition $\mathcal{H}=\bigoplus_{g\in\Gamma}\mathcal{H}_{g}$. The morphisms are linear maps that preserve the $\Gamma$-grading. The tensor product is the usual Hilbert space tensor product with the grading defined by
@@
If $\Gamma$ is a group and $\omega\in Z^{3}(\Gamma,\mathbb{T})$ is a normalised $3$-cocycle, the category $\mathrm{Hilb}(\Gamma,\omega)$ is defined exactly as is $\mathrm{Hilb}(\Gamma)$ but with associators now given by
@@
For the remainder of this paper we will only make use of this example when $\Gamma$ is a group. However, our results also apply to the case of monoids.
@@
Let $A$ be a $*{}^{*}$-algebra. We denote by $\mathrm{Corr}_{0}(A)$ the $*{}^{*}$-tensor category whose objects are non-degenerate $A$-$A$ correspondences and whose morphisms are adjointable intertwiners of the underlying Hilbert bimodules. The tensor product of two $A$-$A$ correspondences $\varphi$ with $\psi$ is given by their composition $\varphi\circ\psi$. The tensor identity of $\mathrm{Corr}_{0}(A)$ is given by the identity homomorphism $\mathrm{id}_{A}$, the associator is given by the rebracketing morphism associated to the underlying tensor product of Hilbert bimodules.
@@
In general, $\mathrm{Corr}(A)$ is not a $*{}^{*}$-tensor category as there is no tensor unit on degenerate correspondences; $A$ does not act as a unit on a degenerate correspondence. It is a non-unital $*{}^{*}$-tensor category, this is a weakening of Definition 1.15 which omits the necessity of a tensor unit.

### 08 Apr 2026 v3

@@ 1. Preliminaries
First, note that $\overline{\psi(B)C}=\psi(B)C$ by Cohen’s factorisation (see [39, Proposition 2.33]). Let $T:{}_{\phi}B\boxtimes{}_{\psi}C\to{}_{\psi\circ\phi}\psi(B)C$ be the continuous linear map given by $T(b\boxtimes c)=\psi(b)c$ for any $b\in B$ and $c\in C$. We claim that $T$ is a bimodule isomorphism. Using the definition of the inner product, a standard check shows that $T$ is a well-defined bimodule map. Taking $(\eta_{\lambda})_{\lambda\in\Lambda}$ to be an approximate unit for $B$ let $S:{}_{\psi\circ\phi}\psi(B)C\to{}_{\phi}B\boxtimes{}_{\psi}C$ be given by $S(c)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes c$ for any $c\in C$. The map $S$ is well-defined precisely because $\psi$ is extendible. Indeed, for $\lambda,\mu\in\Lambda$
@@
Therefore, $S$ is well defined if and only if $\psi(\eta_{\lambda}-\eta_{\mu})c$ converges to 0 for any $c\in C$, which is equivalent to saying that $\psi(\eta_{\lambda})$ converges strictly in $\mathcal{M}(C)$. Hence $S$ is well-defined. Moreover, it can be seen that $S$ is the adjoint of $T$ and hence continuous. Now for any $b\in B$ and $c\in C$,
@@
Let AA be a C-algebra. A Hilbert AA-$B$-bimodule $X$ is called non-degenerate if XBXB is dense in $X$ and AXAX is dense in $X$.
@@
Note that if $X$ is a right-Hilbert $B$-module then XBXB is always dense in $X$ by the argument in [26, Proposition 2.16]. Therefore, a Hilbert AA-$B$-bimodule $E$ may only fail to be non-degenerate if AXAX is not dense in $X$.
We end this subsection by showing that if AA is a $\mathrm{C}^{*}$-algebra and $E$ is a non-degenerate Hilbert AA-bimodule, then we can obtain a well-defined action of $\mathcal{M}(A)$ on $E$.
Let $(\eta_{\lambda})_{\lambda\in\Lambda}$ be an approximate unit of AA and let $L_{E}:E\to A\boxtimes E$ and $R_{E}:E\to E\boxtimes A$ be the AA-bimodule maps given by $L_{E}(x)=\lim\limits_{\lambda}\eta_{\lambda}\boxtimes x$ and $R_{E}(x)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}$ for all $x\in E$. Note that, since $E$ is non-degenerate, $L_{E}$ and $R_{E}$ are unitary bimodule isomorphisms. This follows similarly to the proof of Proposition 1.6. The inverses of $L_{E}$ and $R_{E}$ are given by their adjoints $L_{E}^{-1}(a\boxtimes x)=a\rhd x$ and $R_{E}^{-1}(x\boxtimes a)=x\lhd a$, for all $a\in A$ and $x\in E$ respectively.
@@
Let AA be a $\mathrm{C}^{*}$-algebra and $E$ be a non-degenerate Hilbert AA-bimodule. Then one may extend the left and right actions of AA on $E$ to left and right actions of $\mathcal{M}(A)$ on $E$. These extended actions equip $E$ with the structure of a Hilbert $\mathcal{M}(A)$-bimodule.
@@
[…]
We claim that these formulae define left and right actions of $\mathcal{M}(A)$ on $E$. Moreover, it is clear that (1.1) restricted to AA coincides with the AA-bimodule structure of $E$.
First, note that the left action of $\mathcal{M}(A)$ on $A\boxtimes E$ is given by left multiplication on AA. Similarly, the right action of $\mathcal{M}(A)$ on $E\boxtimes A$ is given by right multiplication on AA. Since $L_{E},R_{E}$, and their inverses are bimodule maps, it is straightforward to see that the formulae in (1.1) define left and right actions of $\mathcal{M}(A)$ on $E$. To see that $E$ with its right AA-valued inner product is a Hilbert $\mathcal{M}(A)$-bimodule it suffices to check that the right $\mathcal{M}(A)$-action commutes with the inner product and the left $\mathcal{M}(A)$-action consists of adjointable operators. First, for $x,y\in E$ and $v\in\mathcal{M}(A)$ we have that
@@
so the right $\mathcal{M}(A)$-action commutes with the right inner product. Moreover, the operator of left multiplication by $v\in\mathcal{M}(A)$ has as an adjoint the operator of left multiplication by $v^{*}$. Indeed left multiplication by $v^{*}$ is the adjoint of the operator of left multiplication by $v$ on the Hilbert AA-module AA so for $x,y\in E$
@@ 1.2. Correspondences
A reformulation of the theory of Hilbert bimodules is the language of C-correspondences.
@@
Let $A,B$ be $\mathrm{C}^{*}$-algebras. An AA-$B$-correspondence is a Hilbert $B$-module $X_{B}$ together with a -homomorphism $\phi:A\to\mathcal{L}(X_{B})$. We often simply denote a correspondence by its underlying -homomorphism. Moreover, we denote the collection of AA-$B$-correspondences by $\mathrm{Corr}(A,B)$.
@@
Note that any AA-$B$-correspondence induces a right-Hilbert AA-$B$-bimodule and vice versa. If $\phi:A\to\mathcal{L}(X_{B})$ is an AA-$B$-correspondence, then $X_{B}$ becomes a right-Hilbert AA-$B$-bimodule, with the left action given by $\phi$. We will often denote this bimodule by $_{\phi}X$, forgetting the right $B$-action. Conversely, given a right-Hilbert AA-$B$-bimodule $X$, the left action by AA induces an AA-$B$-correspondence. Therefore, we will freely flip between the two pictures.
One may compose correspondences through the tensor product of bimodules. Precisely, let $X$ be a Hilbert AA-$B$-bimodule inducing $\phi\in\mathrm{Corr}(A,B)$, and let $E$ be a Hilbert $B$-$C$-bimodule inducing $\psi\in\mathrm{Corr}(B,C)$. Then $X\boxtimes E$ gives a Hilbert AA-$C$-bimodule which induces an element in $\mathrm{Corr}(A,C)$ denoted by $\psi\circ\phi$. Although the theories of bimodules and correspondences are equivalent, we sometimes choose to work with correspondences as the composition resembles composition of -homomorphisms between $\mathrm{C}^{*}$-algebras in a covariant manner.
@@ 1.3. $\mathrm{C}^{*}$-tensor categories
Throughout this section we will assume that the reader is familiar with the standard language of category theory. For a category $\mathcal{C}$ we will use capital letters e.g. $X,Y,$ and $Z$ to denote objects of the category. The space of morphisms between two objects $X,Y\in\mathcal{C}$ will be denoted by $\mathrm{Hom}(X,Y)$. All categories in this section will be $\mathbb{C}$-linear, that is an additive category such that its space of morphisms $\mathrm{Hom}(X,Y)$ between any two objects $X,Y$ is a $\mathbb{C}$-vector space and the composition of morphisms yields a bilinear map $\mathrm{Hom}(X,Y)\times\mathrm{Hom}(Y,Z)\rightarrow\mathrm{Hom}(X,Z)$ for any $X,Y,Z\in\mathcal{C}$.
We invite the reader to recall the definition of a $\mathrm{C}^{*}$-category from [20]. Let $\mathcal{C}$ and $\mathcal{D}$ be $\mathrm{C}^{*}$-categories. A functor $F:\mathcal{C}\rightarrow\mathcal{D}$ is called a $\mathrm{C}^{*}$-functor if the induced mappings $\mathrm{Hom}(X,Y)\to\mathrm{Hom}(F(X),F(Y))$ are $\mathbb{C}$-linear and -preserving. A natural transformation $\nu:F\rightarrow G$ between $\mathrm{C}^{*}$-functors is called an isometry if $\nu_{X}^{*}\nu_{X}=\mathrm{id}_{F(X)}$ for all $X\in\mathcal{C}$. Moreover, $\nu$ is called a unitary if it is a surjective isometry in which case $\nu_{X}\nu_{X}^{*}=\mathrm{id}_{G(X)}$ for all $X\in\mathcal{C}$. We are interested in C-categories that admit tensor product structures.
@@
(see for example [24]) A $\rm{C}^{*}$-tensor category is a $\rm{C}^{*}$-category $\mathcal{C}$ together with a $\mathbb{C}$-linear bifunctor $\otimes-:\mathcal{C}\times\mathcal{C}\rightarrow\mathcal{C}$, a distinguished object $1_{\mathcal{C}}\in\mathcal{C}$ and unitary natural isomorphisms
@@
(1.3) $((W\otimes X)\otimes Y)\otimes Z$$(W\otimes(X\otimes Y))\otimes Z$$(W\otimes X)\otimes(Y\otimes Z)$$W\otimes((X\otimes Y)\otimes Z)$$W\otimes(X\otimes(Y\otimes Z)),$$\scriptstyle{\alpha_{W,X,Y}\otimes\operatorname{id_{Z}}}$$\scriptstyle{\alpha_{W\otimes X,Y,Z}}$$\scriptstyle{\alpha_{W,X\otimes Y,Z}}$$\scriptstyle{\alpha_{W,X,Y\otimes Z}}$$\scriptstyle{\operatorname{id_{W}}\otimes\alpha_{X,Y,Z}}$
(1.4) $(X\otimes 1_{\mathcal{C}})\otimes Y$$X\otimes(1_{\mathcal{C}}\otimes Y)$$X\otimes Y.$$\scriptstyle{\alpha_{X,1,Y}}$$\scriptstyle{\rho_{X}\otimes\mathrm{id}_{Y}}$$\scriptstyle{\mathrm{id}_{X}\otimes\lambda_{Y}}$
and $\mathrm{Hom}(1_{\mathcal{C}},1_{\mathcal{C}})\cong\mathbb{C}$. Moreover, $\mathcal{C}$ is said to be semisimple if there exists a collection of objects $X_{i}\in\mathcal{C}$ with $\mathrm{Hom}(X_{i},X_{j})\cong\delta_{ij}\mathbb{C}$ and any object $X\in\mathcal{C}$ can be decomposed uniquely as a finite direct sum $X\cong\bigoplus_{i}X_{i}^{\oplus m_{i}}$, where $m_{i}$ is the multiplicity of each irreducible $X_{i}$.
We call the structure morphism $\alpha$ as in Definition 1.13 associated to a C-tensor category its associator and the maps $\lambda$ and $\rho$ the unitors. We call an object $X\in\mathcal{C}$ such that $\mathrm{Hom}(X,X)\cong\mathbb{C}$ irreducible or simple. For a C-tensor category $\mathcal{C}$, we will denote by $\mathrm{Irr}(\mathcal{C})$ a collection of isomorphism class representatives for simple objects in $\mathcal{C}$.
@@
Let $\Gamma$ be a countable discrete group (or more generally a countable, discrete monoid). We denote by $\mathrm{Hilb}(\Gamma)$ the semisimple C-tensor category whose objects are finite-dimensional $\Gamma$-graded Hilbert spaces, i.e. finite-dimensional Hilbert spaces $\mathcal{H}$ endowed with a decomposition $\mathcal{H}=\bigoplus_{g\in\Gamma}\mathcal{H}_{g}$. The morphisms are linear maps that preserve the $\Gamma$-grading. The tensor product is the usual Hilbert space tensor product with the grading defined by
@@
If $\Gamma$ is a countable discrete group and $\omega\in Z^{3}(\Gamma,\mathbb{T})$ is a normalised 33-cocycle (see [4, Section III.1] for definitions), the category $\mathrm{Hilb}(\Gamma,\omega)$ is defined exactly as is $\mathrm{Hilb}(\Gamma)$ but with associators now given by
@@
For the remainder of this paper we will only make use of this example when $\Gamma$ is a countable discrete group. However, our results also apply to the case of monoids.
@@
Let AA be a C-algebra. We denote by $\mathrm{Corr}_{0}(A)$ the C-tensor category whose objects are non-degenerate AA-AA-correspondences and whose morphisms are adjointable bimodule maps between the underlying Hilbert bimodules. The tensor product of two AA-AA-correspondences $\varphi$ with $\psi$ is given by their composition $\varphi\circ\psi$. The tensor identity of $\mathrm{Corr}_{0}(A)$ is given by the identity homomorphism $\mathrm{id}_{A}$, the associator is given by the rebracketing morphism associated to the underlying tensor product of Hilbert bimodules.
@@
In general, $\mathrm{Corr}(A)$ is not a C-tensor category as there is no tensor unit on degenerate correspondences; AA does not act as a unit on a degenerate correspondence. It is a non-unital C-tensor category; this is a weakening of Definition 1.13 which omits the necessity of a tensor unit.